### v1

@@ 3.2 One-Step Diffusion (OSD) Model
Model Architecture Overview. Most image restoration tasks with OSD models are extensively studied in previous works Wang et al. (2024e); Wu et al. (2024a); Wang et al. (2024a); Li et al. (2024a). However, these methods struggle to achieve desirable results in human body restoration (HBR). To address this limitation, we propose OSDHuman, an OSD model specifically designed for HBR. Specifically, we adopt a Stable Diffusion (SD) model architecture Rombach et al. (2022) by fixing the number of steps, thereby transforming it into an OSD framework. As shown in Fig. 4, the first step uses the variational autoencoder (VAE) encoder $E_{\theta}$ to project the low-quality (LQ) image $I_{L}$ into the latent space, resulting in $z_{L}=E_{\theta}(I_{L})$. Subsequently, a single denoising operation $F_{\theta}$ is applied to estimate the noise, which is crucial for enabling the calculation of the predicted high-quality (HQ) latent vector $\hat{z}_{H}$ through the equation:
@@
Training Objective. During training, we utilize pixel-wise MSE loss and perceptual loss LPIPS Zhang et al. (2018). Additionally, the obtained $\hat{z}_{H}$ is used to compute the variational score distillation (VSD) loss, ensuring alignment between the generated images and natural images. The final overall training objective for the generator $\mathcal{G}_{\theta}$ is the following. $\lambda_{1}$ and $\lambda_{2}$ are the weights for $\mathcal{L}_{\text{LPIPS}}$ and $\mathcal{L}_{\text{VSD}}$.
@@
High-Fidelity Image Embedder. The degradation level of LQ human images is generally significant, and the image content is often highly complex. It is necessary to consider the coordination between the human pose and the surrendering. Therefore, we propose an HFIE that can guide the restoration direction, reducing the feature gap between HQ and LQ human images. OSEDiff Wu et al. (2024a) employs a finetuned RAM Zhang et al. (2023) as the degradation-aware prompt extractor (DAPE). It provides tags to guide the OSD model. However, the tags generated by DAPE are often too broad and imprecise for human images, offering insufficient and even biased guidance to the OSD model.
@@
As shown in Fig. 5, the RAM used in DAPE consists of two parts: the image encoder and the tagging head. However, HFIE only uses the image encoder ($\mathcal{E}$) in RAM, leveraging the Swin Transformer Liu et al. (2021b), which downsamples the input LQ $I^{\prime}_{L}\in\mathbb{R}^{384\times 384\times 3}$ by a factor of 32. The $I^{\prime}_{L}$ represents the resized LQ $I_{L}$. The image embeddings are $x_{L}=\{x_{L,k}\in\mathbb{R}^{512}\}_{k=1}^{145}$, where the first 144 embeddings represent local information of images. The remaining 1 embedding is obtained through the average pooling of others, which contains the overall information of images. Therefore, using a linear layer to reduce the embeddings’ size to match the input of SD (e.g., 77$\times$1024 for SD-2.1) would result in the loss of distinguish ability between overall and local information. Our proposed method HFIE uses learnable embeddings $Q$ as the query input to the multi-head attention (MHA) layer, ultimately producing the HFIE:
@@
Variational Score Distillation (VSD). Fine-tuning image restoration models often encounter challenges due to the limited training data available, especially when compared to large-scale foundation models like Stable Diffusion (SD). It leads to generated images that fail to align with the natural image space distribution. Previous studies Yin et al. (2023); Wang et al. (2023b); Dao et al. (2024) propose some VSD methods to solve this problem by aligning the distributions represented by two diffusion models. Following OSEDiff Wu et al. (2024a), we apply a VSD module in latent space to guide OSDHuman in learning the distribution of natural images from SD. The VSD loss is computed from the distribution gap in the latent space output by the pretrained regularizer $\epsilon_{\phi}$ and finetuned regularizer $\epsilon_{\phi^{\prime}}$. The gradient of the VSD loss is defined as:
@@
$=\mathop{\mathbb{E}}\limits_{t,\epsilon,\hat{z}_{t}}\left[\frac{\epsilon_{\phi}(\hat{z}_{t};t,p)-\epsilon_{\phi^{\prime}}(\hat{z}_{t};t,p)}{\text{mean}(||\epsilon_{\phi}(\hat{z}_{t};t,p)-\hat{z}_{H}||)}\cdot\frac{\partial\hat{z}_{H}}{\partial\theta}\right],$
@@
Figure 7: The distribution of tags in the PERSONA dataset, identified by the Recognize Anything Plus Model Huang et al. (2023). The angles of the pie chart represent the frequency of each tag category in the dataset, i.e., the tag count. The tag category count indicates how many tags are contained within each category.
@@ 4.1 Quantitative Analysis of PERSONA Dataset
High-Quality Dataset. Our person-based restoration dataset with sophisticated objects and natural activities (PERSONA) dataset, consists of 109,052 human images with a resolution of 512$\times$512. These images are obtained through the high-quality automated cropping and filtering (HQ-ACF) pipeline. To quantify the quality of the PERSONA dataset, we evaluated a range of no-reference image quality assessment (IQA) scores, including CLIPIQA Wang et al. (2023a), MANIQA Yang et al. (2022), MUSIQ Ke et al. (2021), BRISQUE Mittal et al. (2011), and NIQE Zhang et al. (2015). We compare the results with those from other human-related datasets, including the object detection datasets OID Kuznetsova et al. (2020); Krasin et al. (2017), VOC Everingham et al. (2010), COCO Lin et al. (2014), Object365 Shao et al. (2019), CrowdHuman Shao et al. (2018), and fashion domain datasets DeepFashion Liu et al. (2016), iDesigner Dufour et al. (2022). As shown in Tab. 1, our dataset consistently outperforms these datasets across all the no-reference IQA measures.
Rich-Diversity Dataset. We utilize the Recognize Anything Plus Model Huang et al. (2023) to obtain image understanding tags for our PERSONA dataset. The visual results are shown in Fig. 7. In total, 3,365 distinct tag categories are identified from the dataset. And the largest proportion belongs to the “Objects” category, demonstrating the presence of sophisticated objects in our dataset. Additionally, the dataset generates a total of 1,452,088 tags, with approximately half of these tags falling under the “People”, “Sports”, and “Activities” categories. It indicates that the dataset is centered on natural human activities. As shown in Fig. 6, most of the images relate to this theme. Finally, the average number of tags per image in the dataset is 13.32, highlighting the dataset’s ease of understanding.
@@ 4.2 Experimental Settings
Training and Testing dataset. Our OSDHuman model is trained on our PERSONA dataset, which contains 109,052 high-quality 512$\times$512 human images. The degradation pipeline of Real-ESRGAN Wang et al. (2021b) is used to generate synthetic degraded images for training. The test data includes PERSONA-Val and PERSONA-Test, both generated by our HQ-ACF pipeline. The HQ images in the validation set are specially selected from those that comply with the pipeline, ensuring that no images in the validation set share sources with the training set. A total of 4,216 images are used, and the degraded LQ images are generated using the same degradation pipeline as during training. The test set is derived from the VOC dataset Everingham et al. (2010) by performing a partial crop using the HQ-ACF pipeline, followed by sampling under predefined IQA thresholds, yielding 3,000 images with real-world LQ.
Evaluation Metrics. For the PERSONA-Val, we employ both reference-based and non-reference IQA metrics. DISTS Ding et al. (2020) and LPIPS Zhang et al. (2018) are used for reference-based perceptual quality assessment, while PSNR and SSIM Wang et al. (2004) (calculated on the Y channel in YCbCr space) are used for reference-based fidelity assessment. Additionally, FID Heusel et al. (2017) is used to measure the distribution between the restored images and GT. The non-reference IQA metrics we used includes CLIPIQA Wang et al. (2023a), MUSIQ Ke et al. (2021), MANIQA Yang et al. (2022), and NIQE Zhang et al. (2015). For the PERSONA-Test, we use the same non-reference IQA metrics as PERSONA-Val. We utilize the evaluation codes provided by pyiqa Chen & Mo (2022), with the pipal version used for MANIQA.
Implementation Details. The OSDHuman model is trained by AdamW optimizer Loshchilov & Hutter (2019) with $\times 10^{-5}$ learning rate and a batch size of 16. The Stable Diffusion v2-1 model AI (2022) serves as the pretrained OSD model with the timestep frozen to 999, and the prompt embedding is provided by HFIE. The LoRA rank for the VAE encoder, the U-Net of the generator, and the regularizer are all set to 4. The weighting scalars $\lambda_{1}$ and $\lambda_{2}$ in Eq. 4 are set to 2 and 1, respectively. Training is conducted for 35K iterations on 4 NVIDIA A800 GPUs.
Compared Methods. We compare OSDHuman with several diffusion-based methods, including DiffBIR Lin et al. (2024), SeeSR Wu et al. (2024b), PASD Yang et al. (2023), ResShift Yue et al. (2023), SinSR Wang et al. (2024e) and OSEDiff Wu et al. (2024a). Among them, SinSR and OSEDiff are OSD models. ResShift and OSEDiff are retrained on our PERSONA dataset, referred to as ResShift* and OSEDiff*, respectively. Additionally, SinSR is distilled using ResShift*, namely SinSR*.
@@ 4.3 Main Results
Quantitative Comparisons. Table 2 presents the evaluation metrics of our OSDHuman on the synthetic PERSONA-Val and real-world PERSONA-Test datasets. Our method achieves the best or second-best results across most metrics when compared with both original and retrained one-step diffusion models. Against multi-step diffusion methods, OSDHuman outperforms in DISTS, LPIPS, PSNR, SSIM, and NIQE on PERSONA-Val, as well as CLIPIQA, MANIQA, and MUSIQ on PERSONA-Test, with other metrics being comparable. While multi-step diffusion methods excel in reconstructing highly degraded regions and achieving higher CLIPIQA scores, their generated content often deviates from the original, resulting in lower fidelity and perceptual scores such as PSNR, SSIM, DISTS, and LPIPS.
Visual Comparisons. As shown in Figs. 8 and 9, representative images from the real-world PERSONA-Test dataset and the synthetic PERSONA-Val dataset are visualized. Existing state-of-the-art image restoration methods are not well suited for human body restoration (HBR). In HBR tasks, the most challenging aspects often involve parts of the human image with intricate textures and delicate structures, such as faces, fingers, and surrounding objects. These methods frequently exhibit issues like over-smoothing or unnatural color rendering, making it difficult to accurately restore fine details such as facial features. For example, DiffBIR Lin et al. (2024) suffers from over-smoothing or non-faithful textures, OSEDiff Wu et al. (2024a) often results in over-saturated facial regions and ResShift Yue et al. (2023) exhibits distorted facial features. In contrast, our OSDHuman model can restore natural human actions and facial expressions in images, achieving high fidelity and maintaining a high degree of similarity to the original image.
@@ 4.4 Ablation Studies
Comparison on training datasets. To evaluate the suitability of our PERSONA dataset for HBR-specific models, we train our OSDHuman on different datasets. The first option employs LSDIR Li et al. (2023), a dataset widely utilized in the image restoration domain. During training, images in the dataset are randomly cropped to 512$\times$512 as input. The second option employs our PERSONA dataset. As shown in Tab. 3, the results of the model trained on our PERSONA significantly outperform that trained on LSDIR in both PERSONA-Val and PERSONA-Test. This demonstrates that our dataset provides a strong prior for the human body, effectively enhancing the performance of HBR.
Comparison on Prompt Extractors. We conduct experiments with four options to evaluate the effectiveness of various prompt extractors for HBR. The first option does not employ a prompt extractor but uses a space as the prompt. For the second option, we use DAPE Wu et al. (2024b) with the setting of OSEDiff Wu et al. (2024a). In OSEDiff, HQ images are inputted into DAPE during training to provide higher-quality prompts. Following this, we also input HQ images into HFIE as the third option. The last option, which is our default setting, involves using HFIE to extract prompts from LQ images. As shown in Tab. 4, our HFIE of default setting outperforms the other options. It demonstrates the superior capability of HFIE in providing priors to the model, guiding higher-quality restoration.

### v2

@@ 3.2 One-Step Diffusion (OSD) Model
Model Architecture Overview. Most image restoration tasks with OSD models are extensively studied in previous works (Wang et al., 2024d; Wu et al., 2024a; Wang et al., 2024b; Li et al., 2024a). However, these methods struggle to achieve desirable results in human body restoration (HBR). To address this limitation, we propose OSDHuman, an OSD model specifically designed for HBR. Specifically, we adopt a Stable Diffusion (SD) model architecture (Rombach et al., 2022) by fixing the number of steps, thereby transforming it into an OSD framework. As shown in Fig. 4, the first step uses the variational autoencoder (VAE) encoder $E_{\theta}$ to project the low-quality (LQ) image $I_{L}$ into the latent space, resulting in $z_{L}=E_{\theta}(I_{L})$. Subsequently, a single denoising operation $F_{\theta}$ is applied to estimate the noise, which is crucial for enabling the calculation of the predicted high-quality (HQ) latent vector $\hat{z}_{H}$ through the equation:
@@
Training Objective. During training, we utilize pixel-wise MSE loss and perceptual loss LPIPS (Zhang et al., 2018). Additionally, the obtained $\hat{z}_{H}$ is used to compute the variational score distillation (VSD) loss, ensuring alignment between the generated images and natural images. The final overall training objective for the generator $\mathcal{G}_{\theta}$ is the following. $\lambda_{1}$ and $\lambda_{2}$ are the weights for $\mathcal{L}_{\text{LPIPS}}$ and $\mathcal{L}_{\text{VSD}}$.
@@
High-Fidelity Image Embedder. The degradation level of LQ human images is generally significant, and the image content is often highly complex. It is necessary to consider the coordination between the human pose and the surrendering. Therefore, we propose an HFIE that can guide the restoration direction, reducing the feature gap between HQ and LQ human images. OSEDiff (Wu et al., 2024a) employs a finetuned RAM (Zhang et al., 2023) as the degradation-aware prompt extractor (DAPE). It provides tags to guide the OSD model. However, the tags generated by DAPE are often too broad and imprecise for human images, offering insufficient and even biased guidance to the OSD model.
@@
[…]
@@
Variational Score Distillation (VSD). Fine-tuning image restoration models often encounter challenges due to the limited training data available, especially when compared to large-scale foundation models like Stable Diffusion (SD). It leads to generated images that fail to align with the natural image space distribution. Previous studies (Yin et al., 2023; Wang et al., 2023b; Dao et al., 2024) propose some VSD methods to solve this problem by aligning the distributions represented by two diffusion models. Following OSEDiff (Wu et al., 2024a), we apply a VSD module in latent space to guide OSDHuman in learning the distribution of natural images from SD. The VSD loss is computed from the distribution gap in the latent space output by the pretrained regularizer $\epsilon_{\phi}$ and finetuned regularizer $\epsilon_{\phi^{\prime}}$. The gradient of the VSD loss is defined as:
@@
[…]
@@
Figure 7: The distribution of tags in the PERSONA dataset, identified by the Recognize Anything Plus Model (Huang et al., 2023). The angles of the pie chart represent the frequency of each tag category in the dataset, i.e., the tag count. The tag category count indicates how many tags are contained within each category.
@@ 4.1 Quantitative Analysis of PERSONA Dataset
High-Quality Dataset. Our person-based restoration dataset with sophisticated objects and natural activities (PERSONA) dataset, consists of 109,053 human images with a resolution of $\times512$. These images are obtained through the high-quality automated cropping and filtering (HQ-ACF) pipeline. To quantify the quality of the PERSONA dataset, we evaluated a range of no-reference image quality assessment (IQA) scores, including CLIPIQA (Wang et al., 2023a), MANIQA (Yang et al., 2022), MUSIQ (Ke et al., 2021), BRISQUE (Mittal et al., 2011), and NIQE (Zhang et al., 2015). We compare the results with those from other human-related datasets, including the object detection datasets OID (Kuznetsova et al., 2020; Krasin et al., 2017), VOC (Everingham et al., 2010), COCO (Lin et al., 2014), Object365 (Shao et al., 2019), CrowdHuman (Shao et al., 2018), and fashion domain datasets DeepFashion (Liu et al., 2016), iDesigner (Dufour et al., 2022). As shown in Tab. 1, our dataset consistently outperforms these datasets across all the no-reference IQA measures.
Rich-Diversity Dataset. We utilize the Recognize Anything Plus Model (Huang et al., 2023) to obtain image understanding tags for our PERSONA dataset. The visual results are shown in Fig. 7. In total, 3,365 distinct tag categories are identified from the dataset. And the largest proportion belongs to the “Objects” category, demonstrating the presence of sophisticated objects in our dataset. Additionally, the dataset generates a total of 1,452,088 tags, with approximately half of these tags falling under the “People”, “Sports”, and “Activities” categories. It indicates that the dataset is centered on natural human activities. As shown in Fig. 6, most of the images relate to this theme. Finally, the average number of tags per image in the dataset is 13.32, highlighting the dataset’s ease of understanding.
@@ 4.2 Experimental Settings
Training and Testing dataset. Our OSDHuman model is trained on our PERSONA dataset, which contains 109,053 high-quality $\times512$ human images. The degradation pipeline of Real-ESRGAN (Wang et al., 2021b) is used to generate synthetic degraded images for training. The test data includes PERSONA-Val and PERSONA-Test, both generated by our HQ-ACF pipeline. The HQ images in the validation set are specially selected from those that comply with the pipeline, ensuring that no images in the validation set share sources with the training set. A total of 4,216 images are used, and the degraded LQ images are generated using the same degradation pipeline as during training. The test set is derived from the VOC dataset (Everingham et al., 2010) by performing a partial crop using the HQ-ACF pipeline, followed by sampling under predefined IQA thresholds, yielding 3,000 images with real-world LQ.
Evaluation Metrics. For the PERSONA-Val, we employ both reference-based and non-reference IQA metrics. DISTS (Ding et al., 2020) and LPIPS (Zhang et al., 2018) are used for reference-based perceptual quality assessment, while PSNR and SSIM (Wang et al., 2004) (calculated on the Y channel in YCbCr space) are used for reference-based fidelity assessment. Additionally, FID (Heusel et al., 2017) is used to measure the distribution between the restored images and GT. The non-reference IQA metrics we used includes CLIPIQA (Wang et al., 2023a), MUSIQ (Ke et al., 2021), MANIQA (Yang et al., 2022), and NIQE (Zhang et al., 2015). For the PERSONA-Test, we use the same non-reference IQA metrics as PERSONA-Val. We utilize the evaluation codes provided by pyiqa (Chen and Mo, 2022), with the pipal version used for MANIQA.
Implementation Details. The OSDHuman model is trained by AdamW optimizer (Loshchilov and Hutter, 2019) with a batch size of 16 and 5e-5 learning rate. The Stable Diffusion v2-1 model (Stability AI, 2022) serves as the pretrained OSD model with the timestep frozen to 999, and the prompt embedding is provided by HFIE. The LoRA (Hu et al., 2022) rank for the VAE encoder, the U-Net of the generator, and the regularizer are all set to 4. The weighting scalars $\lambda_{1}$ and $\lambda_{2}$ in Eq. 4 are set to 2 and 1, respectively. Training is conducted for 35K iterations on 4 NVIDIA A800 GPUs.
Compared Methods. We compare OSDHuman with several diffusion-based methods, including DiffBIR (Lin et al., 2024), SeeSR (Wu et al., 2024b), PASD (Yang et al., 2024), ResShift (Yue et al., 2023), SinSR (Wang et al., 2024d) and OSEDiff (Wu et al., 2024a). Among them, SinSR and OSEDiff are OSD models. ResShift and OSEDiff are retrained on our PERSONA dataset, referred to as ResShift* and OSEDiff*, respectively. Additionally, SinSR is distilled using ResShift*, namely SinSR*.
@@ 4.3 Main Results
Quantitative Comparisons. Tab. 2 presents the evaluation metrics of our OSDHuman on the synthetic PERSONA-Val and real-world PERSONA-Test datasets. Our method achieves the best or second-best results across most metrics when compared with both original and retrained one-step diffusion models. Against multi-step diffusion methods, OSDHuman outperforms in DISTS, LPIPS, PSNR, SSIM, and NIQE on PERSONA-Val, as well as CLIPIQA, MANIQA, and MUSIQ on PERSONA-Test, with other metrics being comparable. While multi-step diffusion methods excel in reconstructing highly degraded regions and achieving higher CLIPIQA scores, their generated content often deviates from the original, resulting in lower fidelity and perceptual scores such as PSNR, SSIM, DISTS, and LPIPS.
Visual Comparisons. As shown in Figs. 8 and 9, representative images from the real-world PERSONA-Test dataset and the synthetic PERSONA-Val dataset are visualized. Existing state-of-the-art image restoration methods are not well suited for human body restoration (HBR). In HBR tasks, the most challenging aspects often involve parts of the human image with intricate textures and delicate structures, such as faces, fingers, and surrounding objects. These methods frequently exhibit issues like over-smoothing or unnatural color rendering, making it difficult to accurately restore fine details such as facial features. For example, DiffBIR (Lin et al., 2024) suffers from over-smoothing or non-faithful textures, OSEDiff (Wu et al., 2024a) often results in over-saturated facial regions and ResShift (Yue et al., 2023) exhibits distorted facial features. In contrast, our OSDHuman model can restore natural human actions and facial expressions in images, achieving high fidelity and maintaining a high degree of similarity to the original image.
@@ 4.4 Ablation Studies
Comparison on Training Datasets. To evaluate the suitability of our PERSONA dataset for HBR-specific models, we train our OSDHuman on different datasets. The first option employs LSDIR (Li et al., 2023), a dataset widely utilized in the image restoration domain. During training, images in the dataset are randomly cropped to $\times512$ as input. The second option employs our PERSONA dataset. As shown in Tab. 3, the results of the model trained on our PERSONA significantly outperform that trained on LSDIR in both PERSONA-Val and PERSONA-Test. This demonstrates that our dataset provides a strong prior for the human body, effectively enhancing the performance of HBR.
Comparison on Prompt Extractors. We conduct experiments with four options to evaluate the effectiveness of various prompt extractors for HBR. The first option does not employ a prompt extractor but uses a space as the prompt. For the second option, we use DAPE (Wu et al., 2024b) with the setting of OSEDiff (Wu et al., 2024a). In OSEDiff, HQ images are inputted into DAPE during training to provide higher-quality prompts. Following this, we also input HQ images into HFIE as the third option. The last option, which is our default setting, involves using HFIE to extract prompts from LQ images. As shown in Tab. 4, our HFIE of default setting outperforms the other options. It demonstrates the superior capability of HFIE in providing priors to the model, guiding higher-quality restoration.
@@ 5 Conclusion
Impact Statement
This paper presents research aimed at advancing the field of Machine Learning. While there are various potential societal implications of our work, we believe that none of these require particular emphasis here.
Acknowledgments
This work was supported by Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102) and the Fundamental Research Funds for the Central Universities.